OpenStudy (anonymous):

The length of a rectangle is 2x – 3 feet, and the width is 5 feet. What is the area of the rectangle in square feet? A. 7x – 15 B. 10x – 15 C. 10x – 8 D. 7x – 8

OpenStudy (anonymous):

Area = Length * Width A = LW A = (2x-3)*5 A = 5(2x-3) A = ???

OpenStudy (anonymous):

10x-15

OpenStudy (anonymous):

B.)